### v2

@@ III Network and Error Models
We consider protocols which distribute $N$-qubit multipartite GHZ states, defined as |$\ket{\text{GHZ}_{N}}=\frac{1}{\sqrt{2}}(\ket{0}^{N}+\ket{1}^{N})$ over a quantum network. The performance of such multipartite distribution protocols will be sensitive to network parameters, such as the rate and fidelity at which Bell states can be generated. In the following, we define a quantum network model used for protocol performance evaluation.
@@ III-A Network model
The quantum network is modelled as a graph $G=(V,E)$, which is a collection of nodes $v\in V$ connected by a set of edges $e\in E$ such as seen in Fig. 2. Nodes are devices equipped with a single quantum memory allocated for communication, for each adjoining edge. The nodes can store qubits in quantum memories as well as perform single-qubit, two-qubit and measurement operations. We assume that all nodes have identical capabilities. An edge represents a quantum channel which can be used to share a Bell state of the form |$\ket{\phi^{+}}=\frac{1}{\sqrt{2}}(\ket{00}+\ket{11})$ between two adjacent nodes. Such shared Bell states will henceforth be referred to as entanglement links. Entanglement links are generated probabilistically, with probability $p_{e}$ for network edge $e$. A heralded entanglement distribution scheme is assumed, such that a successful entanglement link generation is flagged by a classical signal [54]. Given the single memory per edge assumption, at most one entanglement link can be stored over an edge and entanglement link generation cannot be re-attempted while the memories are still occupied. We also define a link-state graph $G^{\prime}=(V,E^{\prime})$, which describes the location of entanglement links over the quantum network. In $G^{\prime}$, an edge $e=(u,v)$ will be present when nodes $u$ and $v$ share an entanglement link. This definition of the link-state graph means that $G^{\prime}$ will contain all nodes from $G$, even if some do not hold any entanglement links. In such a situation, such nodes would simply be disconnected in $G^{\prime}$. Finally, we model the operation of the network as being performed in discrete timeslots of duration $T_{\text{slot}}$. A timeslot includes the generation of entanglement links, as well as any local operations and classical communication (LOCC) required to generate the GHZ state. We consider the implementation of a quantum memory cutoff. In a discrete timeslot network model, a cutoff $Q_{c}$ defines the maximum number of timeslots an entanglement link can be stored in a quantum memory before being discarded. As the network operates over multiple timeslots, we use the index $t\in\{1,2,\ldots\}$ to denote the $t^{\text{th}}$ timeslot. One timeslot, of period $T_{\text{slot}}$ in the discrete-time network model is sometimes referred to as a round. In this context, $k$ rounds would refer to an ordered set of $k$ sequential timeslots. Finally, as the state of $G^{\prime}$ will vary between timeslots, as entanglement links are generated and discarded, we define a link-state graph $G^{\prime}_{t}$, which represents the state of $G^{\prime}$ during timeslot $t$.
@@ III-B Error model
We consider multipartite state distribution in a network of noisy quantum devices, where imperfect Bell state generation and decoherence are the dominant sources of error. LOCC operations are assumed to be performed instantaneously and ideally which means that any noise in the distributed GHZ state can be attributed to noise from the Bell states used for distribution. Examples of LOCC operations include entanglement swapping and entanglement fusion. We quantify the impact of noise on the distributed states by their fidelity [55]. The fidelity $\mathcal{F}$ is a measure of the overlap between two quantum states with $0\leq\mathcal{F}\leq 1$. For a mixed state $\rho$, such as a state generated under noisy conditions, the fidelity can be defined with respect to a desired pure state |$\ket{\Psi}$ as:
$\mathcal{F}(\Psi,\rho)=\bra{\Psi}\rho\ket{\Psi}$ (1)
@@
$\rho_{w}=w\,\ket{\phi^{+}}\bra{\phi^{+}}+\frac{1-w}{4}\mathbb{I}_{4}$ (3)
A Werner state is a mixed state which is composed of the ideal Bell state with probability $w$, and the maximally mixed state with probability $1-w$. The fidelity of a Werner state with respect to the desired Bell state can be given in terms of the Werner parameter $w$ as $\mathcal{F}=\frac{3w+1}{4}$ [7].
@@ III-C GHZ state generation and fidelity
Fig. 3(b) shows how a 3-qubit GHZ state can be generated by fusing two Bell states [^1] Note that the |$\ket{\phi^{+}}$ Bell state is equivalent to a $2$-qubit GHZ state.. The fidelity of a GHZ state generated by entanglement fusion between two noisy GHZ states will depend on the noise and the choice of qubits that are used for the fusion operation. The fidelity of a GHZ state, $\mathcal{F}_{\text{GHZ}}$, distributed using the star approach, can be expressed as a function of the fidelity of the Bell states $\mathcal{F}_{B}$ to be fused (one per branch $B$) [28]:
@@ IV-A Multipartite state routing
The lower bound follows from the fact that fidelity $\mathcal{F}_{e}$ of a Werner state over edge $e$ is also equal to the probability that it is found in the desired |$\ket{\phi^{+}}$ state [7]. The value of $\mathcal{F}_{LB}$ therefore denotes the probability that all Werner states in $R$ are found in the |$\ket{\phi^{+}}$ state, which would result in the generation of a GHZ state generated with fidelity at least $\mathcal{F}_{LB}$ [43]. We also consider the term $w_{R}$, which is the product of Werner parameters from $R$ as:
@@ IV-D1 Benchmark single-path protocols
         if $e\notin G^{\prime}$ and GenerateLinkSuccess$(e)$ then
@@ IV-D2 Multi-path protocols
[…]
         (_$(\_,v_{c})\leftarrow$ GetRoutingSolution($G,S,\text{type}$)
@@ IV-E Computational complexity
$p$ $0.1$ Entanglement link generation probability: the probability of generating a Bell state between two adjacent nodes over an edge.
@@ V-B Distribution rate and fidelity trade-off
When comparing the star approaches (mp-s and sp-s), we see the mp-s protocol achieved an improvement in distribution rate of up to ×$\times 2.2$ and $16\%$ improvement in fidelity. Further, we see that for both single-path and multi-path routing, the tree approaches (sp-t and mp-t) outperformed the star approaches (sp-s and mp-s). As the tree approaches select the least cost routing solution, the tree approaches, by definition, select routing solutions that give a higher distribution rate and fidelity. This is because the star routing solutions are valid, but not necessarily optimal, tree solutions. Results for different entanglement link generation probabilities $p\in\{0.2,0.3\}$, found in Appendix D, show the same overall trends.
@@ VI Conclusions and further work
We have proposed protocols for distributing multipartite states which use multi-path routing, to select a route dynamically, using knowledge of the location and fidelity of Bell states over the network. The performance of such protocols was compared to single-path approach with fixed routing, considering both the rate and fidelity at which GHZ states are distributed. Results show that the multi-path protocols distributed GHZ states at a higher rate and fidelity compared to single-path routing. The higher fidelity was achieved due to the reduced impact of noise due to decoherence when using multi-path routing. Finally, by selecting an appropriate quantum memory cutoff time, we showed up to an 8.3×8.3$\times$8.3 × increase in distribution rate and up to a 28% improvement in fidelity compared to the single-path routing approach. For the worst-case set of users, the distribution rate achieved by multi-path routing was up to 78 times higher. These results suggest that multi-path routing could be used to improve performance on near-term quantum networks, particularly when qubit operations are noisy.
@@ Appendix A Formal Problem Statement
By choosing edge-weights $c_{e}=-\log(w_{e})$, the Steiner tree selected will maximise $\mathcal{F}_{LB}$. For Bell states, the transformation between $\mathcal{F}$ and $w$ is affine and monotonic ($w\in[0,1]$), so using either $-\log(\mathcal{F}_{e})$ or $-\log(w_{e})$ as edge-costs for the routing algorithms considered will yield the same optimal routing solution.
@@ Appendix D Sensitivity to entanglement link generation probability
The protocols were simulated for entanglement link generation probabilities $p=0.2,0.3$ using otherwise identical assumptions as in Section V-B. However, the results are averaged over 60 sets of users rather than 100. The entanglement link generation probability $p$ was adjusted to $p=0.2$ for Fig. 11(a) and $p=0.3$ for Fig. 11(b). In Fig. 11(b) the results show that when $p=0.2$, largely consistent with the results seen in Section V-B. The multi-path routing protocols provided up to a 9.5×9.5$\times$9.5 × higher distribution rate and $30\%$ higher GHZ state fidelity. Fig. 11(b) shows that for $p=0.3$ the maximum rate and fidelity improvements were ×$\times 7.6$ and $18\%$ respectively. The benefits of multi-path routing reduce as $p$ increases, as all protocols achieve higher distribution rates and fidelities when $p$ is higher. Of the protocols, the highest absolute fidelity will be achieved by the sp-t protocol with $Q_{c}=1$. This uses the least cost routing solution with all entanglement links discarded at the end of each timeslot. Any multi-path routing solution will use at least as many entanglement links, hence reducing the average GHZ state fidelity. However, a cutoff of $Q_{c}=1$ may not be practical in near-term quantum networks. For $Q_{c}>1$, the multi-path protocols can achieve a significant improvement in distribution rate and fidelity.

### v3

@@ III Network and Error Models
[…]
@@ III-A Network model
[…]
@@ III-B Error model
[…]
@@
[…]
A Werner state is a mixed state which is composed of the ideal Bell state with probability $w$, and the maximally mixed state with probability 1w1-w. The fidelity of a Werner state with respect to the desired Bell state can be given in terms of the Werner parameter $w$ as $\mathcal{F}=\frac{3w+1}{4}$ [7].
@@ III-C GHZ state generation and fidelity
[…]
@@ IV-A Multipartite state routing
[…]
@@ IV-D1 Benchmark single-path protocols
    if $e\notin G^{\prime}$ and GenerateLinkSuccess(e)(e) then
@@ IV-D2 Multi-path protocols
  if $\text{type}==\text{star}$ then
    $(\_,v_{c})\leftarrow$ GetRoutingSolution($G,S,\text{type}$)
@@ IV-E Computational complexity
pp $0.1$ Entanglement link generation probability: the probability of generating a Bell state between two adjacent nodes over an edge.
@@ V-B Distribution rate and fidelity trade-off
When comparing the star approaches (mp-s and sp-s), we see the mp-s protocol achieved an improvement in distribution rate of up to $\times 2.2$ and $16\%$ improvement in fidelity. Further, we see that for both single-path and multi-path routing, the tree approaches (sp-t and mp-t) outperformed the star approaches (sp-s and mp-s). As the tree approaches select the least cost routing solution, the tree approaches, by definition, select routing solutions that give a higher distribution rate and fidelity. This is because the star routing solutions are valid, but not necessarily optimal, tree solutions. Results for different entanglement link generation probabilities $p\in\{0.2,0.3\}$, found in Appendix D, show the same overall trends.
@@ VI Conclusions and further work
We have proposed protocols for distributing multipartite states which use multi-path routing, to select a route dynamically, using knowledge of the location and fidelity of Bell states over the network. The performance of such protocols was compared to single-path approach with fixed routing, considering both the rate and fidelity at which GHZ states are distributed. Results show that the multi-path protocols distributed GHZ states at a higher rate and fidelity compared to single-path routing. The higher fidelity was achieved due to the reduced impact of noise due to decoherence when using multi-path routing. Finally, by selecting an appropriate quantum memory cutoff time, we showed up to an $8.3\times$ increase in distribution rate and up to a 28% improvement in fidelity compared to the single-path routing approach. For the worst-case set of users, the distribution rate achieved by multi-path routing was up to 78 times higher. These results suggest that multi-path routing could be used to improve performance on near-term quantum networks, particularly when qubit operations are noisy.
@@ Appendix A Formal Problem Statement
By choosing edge-weights $c_{e}=-\log(w_{e})$, the Steiner tree selected will maximise $\mathcal{F}_{LB}$. For Bell states, the transformation between $\mathcal{F}$ and $w$ is affine and monotonic ($w\in[0,1]$), so using either $\log(\mathcal{F}_{e})$ or $\log(w_{e})$ as edge-costs for the routing algorithms considered will yield the same optimal routing solution.
@@ Appendix D Sensitivity to entanglement link generation probability
The protocols were simulated for entanglement link generation probabilities $p=0.2,0.3$ using otherwise identical assumptions as in Section V-B. However, the results are averaged over 60 sets of users rather than 100. The entanglement link generation probability $p$ was adjusted to $p=0.2$ for Fig. 11(a) and $p=0.3$ for Fig. 11(b). In Fig. 11(a) the results show that when $p=0.2$, largely consistent with the results seen in Section V-B. The multi-path routing protocols provided up to a $9.5\times$ higher distribution rate and $30\%$ higher GHZ state fidelity. Fig. 11(b) shows that for $p=0.3$ the maximum rate and fidelity improvements were $\times 7.6$ and $18\%$ respectively. The benefits of multi-path routing reduce as $p$ increases, as all protocols achieve higher distribution rates and fidelities when $p$ is higher. Of the protocols, the highest absolute fidelity will be achieved by the sp-t protocol with $Q_{c}=1$. This uses the least cost routing solution with all entanglement links discarded at the end of each timeslot. Any multi-path routing solution will use at least as many entanglement links, hence reducing the average GHZ state fidelity. However, a cutoff of $Q_{c}=1$ may not be practical in near-term quantum networks. For $Q_{c}>1$, the multi-path protocols can achieve a significant improvement in distribution rate and fidelity.